OpenStudy (anonymous):

6x - 4(2x + 8) = 16 x = __________ 5x + 4(2x + 2) = -1 x = __________ 8x - 5x = -15 x = __________

OpenStudy (ajprincess):

6x - 4(2x + 8) = 16 First distribute 4(2x + 8) then u will get 6x-4*2x-4*8=16 6x-8x-32=16 -2x-32=16 Can u find x @jennaaanicoleee?

OpenStudy (anonymous):

-24?

OpenStudy (ajprincess):

yup right:)

OpenStudy (anonymous):

2x=16-32

OpenStudy (ajprincess):

5x + 4(2x + 2) = -1 First distribute 4(2x + 2) then u will get 5x+4*2x+4*2=-1 5x+8x+8=-1 13x+8=-1 Can u find x?

OpenStudy (ajprincess):

@Rthy -2x-32=16 -2x=16+32 -2x=48 x=-48/2 =-24

OpenStudy (anonymous):

no...

OpenStudy (ajprincess):

13x+8=-1 13x=-8-1 13x=-9 x=-9/13 Getting it?

OpenStudy (anonymous):

kinda

OpenStudy (ajprincess):

If it confusing plz ask.

OpenStudy (anonymous):

do u divide -9 and 13?

OpenStudy (ajprincess):

 Is it clear or still confusing?